Can you help. if the hypotenuse is 20cm and the other sides are x and 2x, how do I calculate what x is?
Answer by checkley71(8403)

x~2+(2x)~2=20~2 or 5x~2=20~2 or 5x~2=400 or x~2=400/5 or x~2=80 or x=8.94
proof 8.94~2+2*8.94~2=20~2 or 80+320=400 or 400=400